An observer sees 2 spaceship flying apart with speed 0.99c. What is the speed of 1 spaceship viewed by other

1
Expert's answer
2022-03-06T15:16:40-0500

"u'=u\sqrt{1-\frac{v^2}{c^2}}=0.14c."